Find the required rate of return for Security X if its beta is 1.80, if the risk-free rate is .02, and the market return is .09. Find the expected return if the degree of risk aversion decreases by 1.5 percentage points. How did Security X's return change? Why? How much did the market portfolios' return change? How much did Security Y's return change if that Security Y had a beta of 0.5?
Answer #1

Initial return=2%+1.80*(9%-2%)=14.60%

1.
=2%+1.80*(9%-2%-1.5%)=11.9000%
Decrease by 2.7%

2.
=2%+9%-2%-1.5%
=7.5%

3.
Initial return=2%+0.50*(9%-2%)=5.5000%

New return=2%+0.50*(9%-2%-1.5%)=4.7500%
Decrease by 0.75%